OpenStudy (anonymous):

Find the point on the graph f(x)=x^2 that is closest to the point (2,(1/2))

OpenStudy (anonymous):

Using calculus

hartnn (hartnn):

know distance formula ?

OpenStudy (anonymous):

I feel like I should x)

OpenStudy (anonymous):

I dont

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

hartnn (hartnn):

now, one of the point is 2,1/2 can u find other point's co-ordinates in terms of x ??

hartnn (hartnn):

u need to minimize this distance

hartnn (hartnn):

to minimize the distance, u can also minimize the square of distance, that would be much more easier.

OpenStudy (sirm3d):

@hartnn how do make the equation that large?

hartnn (hartnn):

type \large or \huge before the text `\huge x=` will give you \(\huge x=\)

hartnn (hartnn):

@Fey , can u find other point's co-ordinates in terms of x ??

OpenStudy (sirm3d):

oh, so there's huge. i only know large. thnx

OpenStudy (anonymous):

Oh, u mean. a^2+b^2=c^2?

hartnn (hartnn):

not exactly that, but yes.

OpenStudy (anonymous):

I get what u mean. Can u help me do it thhat way?

hartnn (hartnn):

ok, we want a point on f(x)=x^2 let its x co-ordinate be 'x' then its y co-ordinate = ?

OpenStudy (anonymous):

y?

hartnn (hartnn):

and y=f(x)=x^2 so the point is (x,x^2) got this ?

OpenStudy (anonymous):

ohh! Yeah, ok.

hartnn (hartnn):

now u have 2 points, use distance formula to get the distance between them

hartnn (hartnn):

(x,x^2) and (2,1/2)

OpenStudy (anonymous):

ok. gimme a minute

OpenStudy (anonymous):

ok, do I simplify it?

hartnn (hartnn):

as i said, its lot easier to minimize the SQUARE of distance, so put the points in this formula \(\huge d^2={(x_1-x_2)^2+(y_1-y_2)^2}\) First put it and tell me what u get for d^2

OpenStudy (anonymous):

oh, ok. \[d ^{2}=(x-2)^{2}+(x ^{2}-\frac{ 1 }{ 2})^{2}\]

hartnn (hartnn):

absolutely! now to minimize d^2 , what u do ?

OpenStudy (anonymous):

take the derivative?

hartnn (hartnn):

yes! and equate it to ... ?

OpenStudy (anonymous):

0!!! :D

hartnn (hartnn):

yes :) do it and tell me what u get...

OpenStudy (anonymous):

is the derivative of d^2.... 2dd'?

OpenStudy (anonymous):

or just 2d?

hartnn (hartnn):

don't take the derivative of d^2 , name it as f(x)

hartnn (hartnn):

\(f(x)=(x-2)^{2}+(x ^{2}-\frac{ 1 }{ 2})^{2}\)

OpenStudy (anonymous):

ok. I see what u did there. but I don't see why.

hartnn (hartnn):

minimizing distance and square of distance is same thing! u get same answer for both, only that minimizing SQUARE of distance is a lot easier .....

OpenStudy (anonymous):

ok. then. I got \[f'(x)=2(x-2)+2x(x ^{2}-\frac{ 1 }{ 2 })\]

hartnn (hartnn):

2x or 4x ?

OpenStudy (anonymous):

Oops. 4x (/.-)

hartnn (hartnn):

yes, now equate that to 0 and find values of x

OpenStudy (anonymous):

How do I find the x's?

hartnn (hartnn):

\(f'(x)=2(x-2)+4x(x ^{2}-\frac{ 1 }{ 2 })=0\) do algebraic simplification

OpenStudy (anonymous):

1?

hartnn (hartnn):

u get a cubic equation.

hartnn (hartnn):

yeah, 1 is a solution...

OpenStudy (anonymous):

So the answere is (1,1) :D

OpenStudy (sirm3d):

i agree.

hartnn (hartnn):

yep. thats the point

hartnn (hartnn):

i wonder how you got x=1 so fast

OpenStudy (anonymous):

Thanks @hartnn ! without u, I woudlnt have an A in calculus xD

OpenStudy (anonymous):

I just got f(1)

OpenStudy (anonymous):

and that 1^2=1

hartnn (hartnn):

ok.... welcome ^_^